# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ 4.2 Consensus-based C-ColME
Consider a system which evolves according to (10) with $W_{t}=W$ in (11), for $t\geqslant\tilde{\tau}$. With probability at least $21-\delta/$ 2, for $\alpha_{t}=\alpha$, it holds:
$\operatorname*{\mathbb{E}}\left[\lVert{\prescript{}{c}{\hat{\bm{\mu}}}}^{t+1}-{\prescript{}{c}{\bm{\mu}}}\rVert^{4}\right]\in\mathcal{O}\left(\sup_{W_{1},\cdots,W_{\zeta_{D}}}\operatorname*{\mathbb{E}}\left[\lVert{\prescript{}{c}{\hat{\bm{\mu}}}}^{\zeta_{D}}-{\prescript{}{c}{\bm{\mu}}}\rVert^{4}\right]\alpha^{4t}\right)$
$+\mathcal{O}\left(\frac{(1-\alpha)^{4}\cdot\left(1-{1}/{\ln({{\alpha{\lambda_{2,c}}}})}\right)^{2}}{(1-\alpha{\lambda_{2,c}})^{2}}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}\rVert^{4}\right]}{(t+1)^{2}}\right)$
$+\mathcal{O}\left({(1-\alpha)^{2}}\left(1-\frac{1}{\ln{{\alpha}}}\right)^{2}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{(t+1)^{2}}\right),$ (12)
and, for $\alpha_{t}=\frac{t}{t+1}$, it holds:
$+\mathcal{O}\left(\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]\left(\frac{1+\ln t}{1+t}\right)^{2}\right).$ (13)
Proof.
The presence of the memory term and the requirement to consider convergence of the fourth moment (rather than the second) prevent the direct application of existing results for the convergence of the consensus algorithm. The proof begins by decomposing the fourth moment into three terms. The first term is dependent on on $\prescript{}{c}{\hat{\bm{\mu}}}^{\zeta_{D}}-\prescript{}{c}{\bm{\mu}}$, that is the estimate errors at time $\zeta_{D}$. The second term involves the differences $\prescript{}{c}{W}^{t-\tau}({\prescript{}{c}{\bar{\mathbf{x}}}}^{\tau+1}-\prescript{}{c}{P}{\prescript{}{c}{\bar{\mathbf{x}}}}^{\tau+1})$, reflecting the effectiveness of consensus averaging—specifically, how well consensus reduces the gap between local estimates ${\prescript{}{c}{\bar{\mathbf{x}}}}$ and their empirical averages $\prescript{}{c}{P}{\prescript{}{c}{\bar{\mathbf{x}}}}$. This terms is minimized when $\prescript{}{c}{W}=\prescript{}{c}{P}$, which is, as expected, the best consensus matrix. Finally, the third term depends on $\prescript{}{c}{P}{\prescript{}{c}{\bar{\mathbf{x}}}}^{\tau+1}-\prescript{}{c}{\bm{\mu}}$ and represents the minimum error achievable by averaging the estimates of all agents in the component. The remainder of the proof focuses on bounding these three terms using the stochasticity and symmetry of the matrix $\prescript{}{c}{W}$. A recurrent key step in the proof is bounding sums of the kind $\sum_{\tau=t_{0}}^{t}\frac{\beta^{t-\tau}}{\tau+1}$, which we demonstrate to be $\mathcal{O}\left((1-1/\ln\beta)\frac{1}{t+1}\right)$ in Lemma 15. It is then shown that the first two terms are $\mathcal{O}\left(1/t^{4}\right)$, while the final term determines the asymptotic behavior. ∎
Theorem 6 also reveals that the choice of the memory parameter $\alpha_{t}$ affects differently the three terms. In particular, the speedup proportional to the size of the connected component $|\mathcal{CC}_{a}|=n_{c}$, as observed for B-ColME in Theorem 4, is not achieved with a constant $\alpha_{t}=\alpha$. However, it becomes attainable when $\alpha_{t}=\frac{t}{t+1}$, as it is detailed in the subsequent theorem:
Theorem 7.
Consider a graph component $c$ and pick uniformly at random an agent $a$ in $c$. Let $g(x)\coloneqq x\ln^{2}(ex)$ and $\alpha_{t}=\frac{t}{t+1}$. It holds:
@@
$\mathbb{P}\left(\forall t>\tau^{{}^{\prime\prime}}_{a},\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\epsilon\right)\geqslant 1-\delta$
where $\tau_{a}^{{}^{\prime\prime}}=\max\left\{\zeta_{a}^{D},g\!\left(C\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{|\mathcal{CC}_{a}|\epsilon^{4}\delta}\right)\in\tilde{\mathcal{O}}\left(\frac{\tilde{n}_{\frac{\delta}{2}}(\varepsilon)}{|\mathcal{CC}_{a}|}\right)\right\}$.
Proof.
Due to Theorem 3, for $t\geq\zeta_{a}^{D}$, the fourth moment of the estimate error within component $c$ can be bounded as in Theorem 6 w.p. at least $21-\delta/$ 2. Subsequently, we apply Markov inequality to establish a probabilistic bound at a specific time $t$. By employing a union bound over $t$ and suitably majorizing the resultant series, we achieve the final result. ∎
The supplementary material (Sec. D.3) presents convergence results for the case where $\alpha_{t}=\alpha$, a setting that does not enjoy the same speedup factor.
4.3 Choice of the Graph and other Parameters
The selection of the graph $G(\mathcal{A},\mathcal{E})$ plays a pivotal role in the efficacy of our algorithms. We discuss key desirable properties in $G(\mathcal{A},\mathcal{E})$. First, Theorems 4 and 7 show that learning timescales, $\tau_{a}^{\prime}$ and $\tau_{a}^{\prime\prime}$, decrease as the size of collaborating agent groups, $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$, increase. Thus, a highly connected graph is preferred to foster the formation of large class-homogeneous agent groups after the disconnection of inter-class edges. Second, the complexities—both spatial and temporal—of B-ColME and C-ColME are directly proportional to the agent’s degree within the graph. We would like then the degree to be small and possibly uniform across the agents to balance computation across agents. We observe that these two desiderata can be at odds, as a higher degree generally leads to larger groups $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$. A third criterion, specific to B-ColME, is that each agent’s neighborhood should have a tree-like structure extending up to $d$ hops with $d$ as large as possible.
Taking these factors into account, we opt for the class of simple random regular graphs $\mathcal{G}_{0}(N,r)$. These graphs are sampled uniformly at random from the set of all $r$-regular simple graphs with $N$ nodes, i.e., graphs without parallel edges or self-loops, and in which every node has exactly $r$ neighbors (note that an even product $rN$ guarantees the set is not empty). The class $\mathcal{G}_{0}(N,r)$ exhibits strong connectivity properties for small values of $r$. For instance, for any $r\geqslant 3$, the probability that the sampled graph is connected approaches one as $N$ increases. Additionally, the sample graph demonstrates a local tree-like structure with high probability, as discussed in Appendix E. The selection of $r$ illustrates the trade-off mentioned above: a lower $r$ reduces the computational and memory load on each agent, but may also lead to smaller the expected sizes of $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$, affecting the accuracy of estimates.
@@ 4.3 Choice of the Graph and other Parameters
ColME ($r$ comms) $|\mathcal{A}|$ $\frac{1}{\Delta_{a}^{2}}\log\frac{|\mathcal{A}|}{\Delta_{a}\delta}+\frac{|\mathcal{A}|}{r}+\frac{1}{|\mathcal{C}_{a}|}{\frac{1}{\varepsilon^{2}}\log\frac{1}{\delta\varepsilon^{2}}}$ $\frac{1}{\Delta_{a}^{4}}\frac{|\mathcal{A}|}{\delta}+\frac{|\mathcal{A}|}{r}+\frac{1}{|\mathcal{C}_{a}|}\frac{1}{\delta\varepsilon^{4}}$
B-ColME $rd$ $\frac{1}{\Delta_{a}^{2}}\log\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|\mathcal{CC}_{a}|r}}{\Delta_{a}\delta}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}+\frac{1}{|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{CC}_{a}^{d}}|}{\frac{1}{\varepsilon^{2}}\log\frac{1}{\delta\varepsilon^{2}}}$ $\frac{1}{\Delta_{a}^{4}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|\mathcal{CC}_{a}|r}}{\delta}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}+\frac{1}{|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{CC}_{a}^{d}}|}\frac{1}{\delta\varepsilon^{4}}$
C-ColME ($\alpha_{t}=\frac{t}{t+1}$) $r$- $\frac{1}{\Delta_{a}^{4}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|\mathcal{CC}_{a}|r}}{\delta}+\frac{1}{|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{CC}_{a}}|}\frac{1}{\delta\varepsilon^{4}}$
@@
A final key parameter for B-ColME is the maximum distance $d$ over which local estimates from agents are propagated. This parameter must be carefully calibrated: it should be small enough to ensure that $\mathcal{CC}_{a}^{d}$, for a randomly chosen $a\in\mathcal{A}$, has a tree structure with high probability. However, choosing a $d$ that is too small could unnecessarily restrict the size of $\mathcal{CC}_{a}^{d}$, thereby undermining the effectiveness of the estimation process (Theorem 4). For a comprehensive analysis on how the parameters $r$ and $d$ influence both the structure of $\mathcal{N}_{a}^{d}$ and the size of $\mathcal{CC}_{a}^{d}$, we direct the reader to Appendix E. In this section, we informally summarize the main result:
@@
By selecting $d=\Big{\lfloor}\frac{1}{2}\log_{r-1}\frac{|\mathcal{A}|}{\log_{r-1}|\mathcal{|}A|}\Big{\rfloor}$ the number of nodes $a\in\mathcal{A}$, whose $d$-neighborhood is not a tree, is $o(|\mathcal{A}|)$ with a probability tending to 1 as $|\mathcal{A}|$ increases. For $r\in\Theta\left(\log(1/\delta)\right)$, $|\mathcal{CC}_{a}^{d}|$ (with $d$ chosen as above) is in $\Omega(|\mathcal{A}|^{\frac{1}{2}-\phi})$ for any arbitrarilyy small $\phi>0$ with probability arbitrarily close to 1.
Finally, for C-ColME, the consensus matrix $W$ could be chosen to minimize the second largest module $\lambda_{2,c}$ of the eigenvalues of each block $\prescript{}{c}{W}$ in order to minimize the bound in Theorem 6. This optimization problem has been studied by Xiao and Boyd [2004] and requires in general a centralized solution. In what follows, we consider the following simple, decentralized configuration rule: $(W_{t})_{a,a^{\prime}}=(\max\{|\mathcal{C}_{a}^{t}|,|\mathcal{C}_{a}^{t}|\}+1)^{-1}$ for all $a^{\prime}\in\mathcal{C}_{a}^{t}$ and $(W_{t})_{a,a}=1-\sum_{a^{\prime}\in\mathcal{C}_{a}^{t}}(W_{t})_{a,a^{\prime}}$.
@@ 5 Algorithms’ Comparison
Table 1 presents a comparative analysis of the three algorithms: ColME, B-ColME, and C-ColME. For a fair comparison, we consider a variant of ColME, where each agent is allowed to communicate with $r$ agents at each time $t$. This adjustment guarantees that all three algorithms incur the same communication overhead.
@@
The second column in Table 1 details the space and time complexities for each algorithm. Remarkably, even when $r$ and $d$ are allowed to increase logarithmically with the number of agents $|\mathcal{A}|$, B-ColME maintains its efficiency advantage over ColME. Moreover, C-ColME demonstrates further improvements, reducing the per-agent burden beyond the savings achieved by B-ColME.
@@
The third and fourth columns detail the characteristic times required to achieve $(\epsilon,\delta)$ convergence for the estimates generated by the three algorithms when local data distributions are sub-Gaussian and when they simply exhibit bounded fourth moments. The characteristic times correspond to $\tau_{a}$, $\tau_{a}^{\prime}$, and $\tau_{a}^{\prime\prime}$ in Theorems 24, and 7, respectively. The table reports their asymptotic behavior as the number of agents $|\mathcal{A}|$ increases ignoring logarithmic factors. The detailed derivations and analyses that underpin these results are provided in Appendix F.
@@
Three factors contribute to the characteristic times. The first factor corresponds to the time required to correctly identify the potential collaborators. In the case of ColME, this involves each agent classifying the other $|\mathcal{A}|-1$ agents, leading to a term that scales as $\log|\mathcal{A}|$ or $|\mathcal{A}|$, contingent upon the assumed properties of the local distribution. For B-ColME and C-ColME, $|\mathcal{A}|$ is replaced by $|\mathcal{CC}_{a}|r$, representing an upper bound on the number of connections agents in $\mathcal{CC}_{a}$ may have initially established with agents from different classes. This substitution is not immediately intuitive, as one might initially anticipate the relevant scale to be simply $r$. However, this adjustment accounts for the potential ripple effect of classification errors: a mistake by any agent $a$ can adversely affect the estimates of all agents within the same connected component $\mathcal{CC}_{a}$.
@@
The second factor contributing to the characteristic times is the time each agent needs to collect all relevant information. In the case of ColME, this time is proportional to $|\mathcal{A}|/r$, as an agent queries all other agents. For B-ColME, the time is specifically tied to $d$, the maximum number of hops messages propagate. We observe that the corresponding term does not appear for C-ColME, because it is dominated by the final term.
@@
The final term pertains to the period necessary for achieving accurate mean estimation once the collaborators are identified and showcases the benefits of collaboration. In ColME, the collaboration’s benefit is particularly striking, as all agents within the same class unite to improve their estimates. This collective effort effectively reduces the convergence time by a factor proportional to the size of the collaborating group, $|\mathcal{C}_{a}|$. For B-ColME and C-ColME, although the speed-up remains proportional to the number of collaborating agents, the actual numbers of collaborators, $|\mathcal{CC}_{a}^{d}|$ for B-ColME and $|\mathcal{CC}_{a}|$ for C-ColME, are in general smaller.
@@
In conclusion, while ColME potentially offers the most accurate estimates, it does so with longer convergence times and greater demands on memory and computation for each agent. In contrast, B-ColME and C-ColME present more efficient alternatives, achieving faster convergence with reduced per-agent resource requirements. However, this efficiency may come at the expense of the maximum attainable accuracy. In the next section, we quantify this trade-off experimentally.
@@
Figure 1: Fraction of agents whose estimates deviate from the true value by more than $\epsilon$ (top) and a fraction of neighbors that have not yet been identified as belonging to a different class (bottom) for B-ColME (left) and C-ColME (right). Averages and 95% confidence intervals were computed over 20 realizations.

### v4

@@ 4.2 Message-passing Algorithm: B-ColME
[Proof in Appendix E] Provided that $\mathcal{CC}^{d}_{a}$ is a tree, for any $\delta\in(0,1)$, employing B-ColME, we have:
$\mathbb{P}\left(\forall t>\tau^{B}_{a},\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geqslant 1-\delta$
where $\tau^{B}_{a}=\max\left[\zeta_{a}^{D}+d,\frac{\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)}{|\mathcal{CC}^{d}_{a}|}+d\right]$ and $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)=\Big{\lceil}-\frac{2\sigma^{2}}{\varepsilon^{2}}\ln\left(\frac{\delta}{4}(1-\mathrm{e}^{-\frac{\varepsilon^{2}}{\sigma^{2}}})\right)\Big{\rceil}$ for SGD and as in Theorem 3 for BFMD.
Similar considerations to those for Theorem 3 apply. The additional term $d$ accounts for the delay introduced by the message-passing scheme.
Corollary 5.
@@
$\mathbb{P}\left(\forall t>\tau^{B}_{a}\,:\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geqslant 1-\delta-\delta^{\prime}.$
Theorem 23 (Appendix G) provides upper bounds for $\delta^{\prime}$ when $\mathcal{G}$ is a random regular graph. In particular, as long we set $d$ as in Proposition 6 (Sec. 4.3), $\delta^{\prime}$ converges to 00 as the number of agents $|\mathcal{A}|$ increases.
@@
ColME ($r$ communications) $|\mathcal{A}|$ $\frac{1}{\Delta_{a}^{2}}\log\frac{|\mathcal{A}|}{\Delta_{a}\delta}+\frac{|\mathcal{A}|}{r}+\frac{1}{|\mathcal{C}_{a}|}{\frac{1}{\varepsilon^{2}}\log\frac{1}{\delta\varepsilon^{2}}}$ $\frac{1}{\Delta_{a}^{4}}\frac{|\mathcal{A}|}{\delta}+\frac{|\mathcal{A}|}{r}+\frac{1}{|\mathcal{C}_{a}|}\frac{1}{\delta\varepsilon^{4}}$
C-ColME [Thm. 3] ($\alpha_{t}=\frac{t}{t+1}$) $r$- $\frac{1}{\Delta_{a}^{4}}\frac{|\mathcal{CC}_{a}|r}{\delta}+\frac{1}{|\mathcal{CC}_{a}|}\frac{1}{\delta\varepsilon^{4}}$
B-ColME [Thm. 4] $rd$ $\frac{1}{\Delta_{a}^{2}}\log\frac{|\mathcal{CC}_{a}|r}{\Delta_{a}\delta}+d+\frac{1}{|\mathcal{CC}_{a}^{d}|}{\frac{1}{\varepsilon^{2}}\log\frac{1}{\delta\varepsilon^{2}}}$ $\frac{1}{\Delta_{a}^{4}}\frac{|\mathcal{CC}_{a}|r}{\delta}+d+\frac{1}{|\mathcal{CC}_{a}^{d}|}\frac{1}{\delta\varepsilon^{4}}$
@@
4.3 Choice of the Graph and other Parameters
The selection of the graph $\mathcal{G}(\mathcal{A},\mathcal{E})$ is crucial for the effectiveness of our algorithms. Here we state the key desirable properties of $\mathcal{G}(\mathcal{A},\mathcal{E})$. First, Theorems 3 and 4 show that learning timescales, $\tau^{B}_{a}$ and $\tau^{C}_{a}$, decrease as the size of the collaborating agent groups, $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$, increase. Therefore, a highly connected graph is preferred to promote the formation of large clusters of agents belonging to the same similarity class after the disconnection of inter-class edges. Second, the spatial and temporal complexities of B-ColME and C-ColME are directly proportional to the agents’ degree within the graph. Hence, we want the degree to be small and possibly uniform across the agents to balance computation across agents. Note that the first two criteria partially conflict, as a higher degree generally leads to larger groups $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$, while a smaller degree ensures better spatial and temporal complexities. A third criterion, specific to B-ColME, is that each agent’s neighborhood should have a tree-like structure extending up to $d$ hops, with $d$ as large as possible.
Considering these criteria, we opt for the class of simple random regular graphs $\mathcal{G}_{0}(N,r)$. These graphs are sampled uniformly at random from the set of all $r$-regular simple graphs with $N$ nodes, i.e., graphs without parallel edges or self-loops, and in which every node has exactly $r$ neighbors. Note that an even product $rN$ guarantees the set is not empty. The class $\mathcal{G}_{0}(N,r)$ exhibits strong connectivity properties for small values of $r$. Specifically, for any $r\geqslant 3$, the probability that the sampled graph is connected approaches one as $N$ increases. Moreover, the sampled graph demonstrates a local tree-like structure with high probability (proof in Appendix G). The choice of $r$ (agents’ degree) illustrates the trade-off discussed above between reducing complexity (low $r$) and having large connected components (high $r$). A sensible rule is to select $r$ sufficiently large to guarantee that most agents in the smallest (most critical) class belong to the same connected component. Consider a class including a fraction $p_{k_{a}}$ of agents, Table 3 in Appendix G shows the average fraction of agents in this class that is not connected to the main connected component as a function of $r$. To keep this fraction below e.g. $10^{-2}$ a good rule of thumb is $r=4/p_{k_{a}}$.
A final key parameter for B-ColME is the maximum distance $d$ over which local estimates from agents are propagated. This parameter must be carefully calibrated: it should be small enough to ensure that $\mathcal{CC}_{a}^{d}$, for a randomly chosen $a\in\mathcal{A}$, has a tree-like structure with high probability. However, choosing a $d$ that is too small could unnecessarily restrict the size of $\mathcal{CC}_{a}^{d}$, thereby undermining the effectiveness of the estimation process (Theorem 4). A comprehensive analysis of how the parameters $r$ and $d$ influence both the structure of $\mathcal{N}_{a}^{d}$ and the size of $\mathcal{CC}_{a}^{d}$ can be found in Appendix G. Here, we informally summarize the main result:
@@ 4.3 Choice of the Graph and other Parameters
By selecting $d=\Big{\lfloor}\frac{1}{2}\log_{r-1}\frac{|\mathcal{A}|}{\log_{r-1}|\mathcal{|}A|}\Big{\rfloor}$ the number of nodes $a\in\mathcal{A}$, whose $d$-neighborhood is not a tree, is $o(|\mathcal{A}|)$ with a probability tending to 1 as $|\mathcal{A}|$ increases. For the same $d$ and $r\in\Theta\left(\log(1/\delta)\right)$, $|\mathcal{CC}_{a}^{d}|$ is in $\Omega(|\mathcal{A}|^{\frac{1}{2}-\phi})$ for any arbitrarily small $\phi>0$ with probability arbitrarily close to 1.
Finally, for C-ColME, the consensus matrix $W$ could be chosen to minimize the second largest module $\lambda_{2,c}$ of the eigenvalues of each block $\prescript{}{c}{W}$ in order to minimize the bound in Theorem 2. This optimization problem has been studied by Xiao and Boyd (2004) and requires in general a centralized solution. In what follows, we consider the following simple, decentralized configuration rule: $(W_{t})_{a,b}=\frac{1}{\max\{|\mathcal{C}_{a}^{t}|,|\mathcal{C}_{b}^{t}|\}+1},\forall b\in\mathcal{C}_{a}^{t}$ and $(W_{t})_{a,a}=1-\sum\limits_{b\in\mathcal{C}_{a}^{t}}(W_{t})_{a,b}$, making $W$ symmetric and doubly stochastic.
@@ 5 Algorithms’ Comparison
Table 2 presents a comparative analysis of the three algorithms: ColME, C-ColME, and B-ColME. For a fair comparison, we consider a variant of ColME, where each agent can communicate with $r$ agents at each time $t$, so that all three algorithms incur the same communication overhead.
@@
The second column of Table 2 outlines the space and time complexities of the algorithms. Notably, even when $r$ and $d$ are allowed to increase logarithmically with the number of agents $|\mathcal{A}|$, B-ColME retains its efficiency advantage over ColME. C-ColME demonstrates even greater improvements, further reducing the per-agent burden compared to the savings achieved by B-ColME.
@@
The third and fourth columns detail the characteristic times required to achieve $(\epsilon,\delta)$ convergence for the estimates generated by the three algorithms, considering both sub-Gaussian local data distributions and distributions with bounded fourth moment. The characteristic times correspond to $\tau_{a}$, $\tau^{C}_{a}$, and $\tau^{B}_{a}$ in Theorem 7 in Appendix A, Theorem 3, and 4, respectively. The table reports their asymptotic behavior as the number of agents $|\mathcal{A}|$ increases ignoring logarithmic factors. The detailed derivations of these results are provided in Appendix H.
@@
Three factors contribute to the characteristic times. The first factor is the time required to correctly identify potential collaborators. For ColME, this involves each agent classifying the other $|\mathcal{A}|-1$ agents, leading to a term that scales as $\log|\mathcal{A}|$ or $|\mathcal{A}|$, depending on the assumed properties of the local distribution. For B-ColME and C-ColME, $|\mathcal{A}|$ is replaced by $|\mathcal{CC}_{a}|r$, which represents an upper bound on the number of connections the agents in $\mathcal{CC}_{a}$ may have initially established with agents from different classes. This substitution may not be immediate, as one might initially expect the relevant scale to be simply $r$. However, this adjustment accounts for the potential ripple effect of classification errors: a mistake by any agent $a$ can impact the estimates of all agents within the same connected component $\mathcal{CC}_{a}$.
@@
The second factor contributing to the characteristic times is the time each agent needs to collect all relevant information. For ColME, this time is proportional to $|\mathcal{A}|/r$, as an agent queries all other agents. For B-ColME, the time is specifically tied to $d$, the maximum number of hops messages propagate. Notably, this term does not appear for C-ColME, as it is dominated by the final term.
@@
The final term represents the time needed for accurate mean estimation after collaborators have been identified, highlighting the benefits of collaboration. In ColME, the collaboration’s benefit is particularly striking, as all agents within the same class work together to improve their estimates. This collective effort effectively reduces the convergence time by a factor proportional to the size of the collaborating group, $|\mathcal{C}_{a}|$. For B-ColME and C-ColME, although the speed-up remains proportional to the number of collaborating agents, the actual numbers of collaborators, $|\mathcal{CC}_{a}^{d}|$ for B-ColME and $|\mathcal{CC}_{a}|$ for C-ColME, are in general smaller.
@@
In conclusion, while ColME potentially offers the most accurate estimates, it requires longer convergence times and greater memory and computational resources. In contrast, B-ColME and C-ColME present more efficient alternatives, achieving faster convergence with reduced per-agent resource demands. However, this efficiency may come at the expense of the maximum attainable accuracy. The next section quantifies this trade-off experimentally.
@@
Figure 1: Fraction of agents with estimate deviates by more than $\epsilon$ from the true value, i.e., |$|\{a\in\mathcal{A}:|\hat{\mu}_{a}^{t}-\mu_{a}|>\varepsilon\}|/|\mathcal{A}|$ (top) and fraction of wrong links (bottom) for B-ColME (a) and C-ColME (b), over 20 realizations with 95% confidence intervals.